Select any country worldwide and search out its population and growth rate in 2010. Then use the equation bellow to predict the countries population in 2020, 2030, 2040, 2050, and 2060. Please show your work. You may construct a table similar to the one below to demonstrate your answers. (Note: You may choose any set of 5 ten year increments if you have difficulty finding 2010 data.) Use N=N0ekt

N = The population in the year you are trying to predict
N0 = Starting population
e = mathematical constant = 2.718
k = growth rate expressed as a decimal. For instance a growth rate of 2% has a k=0.02
t = time in future years

Country _______________ Growth Rate ________________ Starting Population ____________ in year __________.

Year

Work

Predicted Population

2010

2020

2030

2040

2050

2060
Answer #1

country -- Australia

growth rate --- 1.6% annually

starting population = 22 million

22 ( 1 + .016 )^t = 22 e^( kt )

1.016 ^t = e^kt

k = ln 1.1016

k = 0.01587

y = 22 e^( 0.01587 t )

year work predicted population
2010 22 e^( 0.01587 *0 ) 22 million
2020 22 e^( 0.01587 *10) 26 million
2030 22 e^( 0.01587 *20 ) 30 million
2040 22 e^( 0.01587 *30 ) 35 million
2050 22 e^( 0.01587 *40) 42 million
2060 22 e^( 0.01587 *50) 49 million